# Preparation of high fidelity entangled cat states with composite pulses

### v1

@@ IV Error and decoherence analysis
In this section, we will investigate the performance of the optimized protocol. Specifically, we discuss the effects of parameters errors on the fidelity of entangled cat state in the Sec. IV.1. Then we analyze the robustness of the entangled cat state to dephasing and single-photon losses in Sec. IV.2. Here, we choose the values $N\in[1,4,8]$.
@@ IV.1 Robustness against parametric errors
It is inevitable that some parametric errors will cause infidelity in the entangled cat states because of the imperfect experimental operations. In this part, with the help of the numerical simulation, we will show the sensitivity of the protocol against parameter errors i.e., timing error $\delta_{t}$, detuning error $\delta_{\Delta}$, and coupling error $\delta_{J}$.
In Fig. 5, we numerically study the robustness of the protocol for timing error $\delta_{t}$. The error range is $\delta_{t}\in[-0.1,0.1]$ and the curve for $N=1$ corresponds to the unoptimized one. As expected, when $N$ takes different value, the fidelity of the entangled cat state changes. Comparing the curves with $N=1$ and $N=4$, we can see the fidelity has been greatly improved. When $N$ is set to a larger value i.e., $N=8$, the fidelity of entangled cat state is almost unaffected by timing error. That is, the protocol can enhance the robustness against timing error.
In actual experiments, imperfect experimental operations may cause inevitable detuning errors. We consider the range of detuning error $\delta_{\Delta}$ as $[-0.1,0.1]$, and plot the fidelity versus detuning error $\delta_{\Delta}$ in the Fig. 5. It can be seen from the figure that the fidelity of the entangled cat state for $N>1$ is much higher than that with $N=1$. The inset in Fig. 5 shows that when the error range is $[-0.05,0.05]$, the fidelities are large than 0.99 when $N>1$ This indicates that the increase of $N$ can reduce the sensitivity to detuning errors.
Figure 5 shows the fidelity against coupling error $\delta_{J}$. As shown in Fig. 5, the fidelity is unaffected by the values of $N$, and the value of fidelity always greater than 0.97. That shows that the protocol cannot improve the robustness against errors in the coupling.
@@ IV.2 Robustness against decoherence
Figure 6: (a) Fidelity versus the single-photon losses rate $\kappa\in[0,0.1]$ $\mathrm{MHz}$ and the dephasing rate $\gamma=0$. (b) Fidelity versus dephasing $\gamma\in[0,0.1]$ $\mathrm{MHz}$ with single-photon losses rate $\kappa=0$.
In the process of gate operation, decoherence can cause infidelity. For the system discussed, we consider two types of noise: single-photon losses and pure dephasing. The master equation of the system can be written as:
@@
where $\rho$ is the density operator of the system, $D[o]\rho=o\rho o^{{\dagger}}-\frac{1}{2}(o^{{\dagger}}o\rho+\rho o^{{\dagger}}o)$. For simplicity, we assume that $\kappa_{k}=\kappa$ $(k=1,2)$ is the single-photon losses rate, and $\gamma_{k}=\gamma$ is the pure dephasing rate. Note that the influence of decoherence in the cavity mode $\mathcal{A}_{0}$ is different from those in the KPOs. The decoherence of the cavity mode $\mathcal{A}_{0}$ can be adiabatically eliminated for large $\Delta$ Chen et al. (2022), so the system is insensitive to the decoherence of $\mathcal{A}_{0}$. In the numerical simulation, we set $\kappa_{0}=\kappa$ and $\gamma_{0}=\gamma$ for convenience. To well understand the effect of decoherence, we project the system onto the eigenstates of $H_{k}^{\mathrm{Kerr}}$. Here, the projection operator of the KPOs is
@@
Figure 7: Fidelity of the optimized protocol with $N=8$ versus $\kappa=\gamma\in[0,0.05]$ $\mathrm{MHz}$. Other parameters remain unchanged.
@@
When $\gamma_{k}$, $\kappa_{k}$ are much smaller than the energy gap $E_{\mathrm{gap}}$, the dynamics of the system is still well confined to the cat-subspace Cai et al. (2021); Puri et al. (2019, 2017, 2020). For large $\alpha$, the influence of the single-photon losses in the KPOs is described by the penultimate term in Eq. (47)
@@
$D^{\prime}[a_{k}]\rho\simeq\frac{\alpha^{2}}{\sqrt{1-e^{-4\alpha}}}D[\sigma^{x}_{k}+ie^{-2\alpha^{2}}\sigma_{k}^{y}]\rho,$ (57)
where $\sigma^{x}_{k}=|C_{+}\rangle_{k}\langle C_{-}|+|C_{-}\rangle_{k}\langle C_{+}|$, $\sigma^{y}_{k}=i(|C_{-}\rangle_{k}\langle C_{+}|+|C_{+}\rangle_{k}\langle C_{-}|)$. Bringing the Eq. (57) into the Eq. (47), we find that in the computational subspace, single-photon losses mainly lead to a bit-flip error $\sigma^{x}_{k}$, accompanied by an exponentially small phase flip error $\sigma^{y}_{k}$.
In Fig. 6, we plot the fidelity of the single-photon losses for different values of $N$. The single-photon losses rate is assumed as $\kappa\in[0,0.1]$ $\mathrm{MHZ}$, and the dephasing rate is assumed as $\gamma=0$. As can be seen from Fig. 6, the increase of $N$ does not affect the fidelity too much. The fidelities can all reach above 0.9 when $\kappa=0.1\mathrm{MHZ}$, which means that the optimized protocol still maintains a high fidelity in the presence of the single-photon losses. We also numerically simulate the dependence of the fidelity on the pure dephasing rate in Fig. 6 and a similar result can be obtained.
Since single-photon losses and dephasing coexist in the system, so we numerically investigate the effect on the fidelity when single-photon losses and dephasing are present together. In Fig. 7, we plot the fidelity of the optimized protocol with $N=8$ versus $\kappa=\gamma\in[0,0.05]$ $\mathrm{MHz}$. The result shows that the optimized protocol maintains a high fidelity more than 0.9 even when photon losses and dephasing exist simultaneously.
@@ V CONCLUSION
In conclusion, we have investigated a protocol to use photonic cat-state qubits for preparing high-fidelity entangled cat states. The entangled cat states can be successfully obtained under the action of the cat-code MS gate. We further adopt composite pulses to improve the fidelity of the entangled cat states. The modulated time-independent composite drives can be realized by introducing composite two-photon squeezing drives. Numerical simulation results show that the protocol can provide strong robustness against timing error. The protocol also exhibits high insensitivity to other parametric errors with proper parameters. Furthermore, under the influence of decoherence, the protocol still maintains a high fidelity. That is, the higher fidelity of the entangled cat states are possible with the larger $N$ of the parameter. On the other hand, employing composite pulses to prepare entangled states is meaningful to investigate. We hope that the proposal could offer a simple and feasible method for preparing stable and high-fidelity entangled cat states in quantum computation.
@@ Acknowledgements.
Y. Xia was supported by the National Natural Science Foundation of China under Grants Nos. 11575045, 11874114, the Natural Science Funds for Distinguished Young Scholar of Fujian Province under Grant 2020J06011 and Project from Fuzhou University under Grant JG202001-2. Y.-H. Chen was supported by the National Natural Science Foundation of China under Grants No. 12304390.

### v2

@@ IV Error and decoherence analysis
In this section, we will investigate the performance of the optimized protocol. Specifically, we discuss the effects of parameters errors on the fidelity of entangled cat state in the Sec. IV.1. Then we analyze the robustness of the protocol to dephasing and single-photon losses in Sec. IV.2. Here, we choose the values $N\in[1,2,4,8]$.
@@ IV.1 Robustness against parametric errors
It is inevitable that imperfect experimental operations will cause infidelity in the entangled cat states. In this part, with the help of the numerical simulation, we will show the sensitivity of the protocol against parameter errors i.e., timing error $\delta_{t}$, detuning error $\delta_{\Delta}$, and coupling error $\delta_{J}$.
In Fig. 6, we numerically study the robustness of the protocol for timing error $\delta_{t}$. The error range is $\delta_{t}\in[-0.1,0.1]$ and the curve for $N=1$ corresponds to the unoptimized one. As it is shown, when $N=2$, the fidelities of the entangled cat states against timing errors have been greatly improved. Furthermore, when $N$ is set to larger values i.e., $N=4$ or $N=8$, the fidelities are further enhanced and are almost unaffected by timing errors. That is, with the increase of $N$, the robustnesses of the protocol to timing errors are enhanced.
In actual experiments, imperfect experimental operations may cause inevitable detuning errors. We consider the range of detuning error $\delta_{\Delta}$ as $[-0.1,0.1]$, and plot the fidelity versus detuning error $\delta_{\Delta}$ in the Fig. 6. It can be seen from the figure that the fidelities of the entangled cat state for $N=2$ are much higher than that with $N=1$ when detuning errors exist. When $N$ increases, the fidelities become higher. The inset in Fig. 6 shows that when the error range is $[-0.05,0.05]$, the fidelities are large than 0.99 when $N>1$ This indicates that the increase of $N$ can reduce the sensitivity to detuning error.
Figure 6 shows the robustness against coupling error $\delta_{J}$. As shown in the figure, the variation of the coupling has a relatively small influence on the fidelities, and the fidelities against coupling errors are independent of the number of $N$. The reason can be seen from Eq. (32), where the coupling strength $J$ is separated from the summation term of $N$. Hence the fidelities of the entangled cat states against coupling errors are also independent of $N$.
@@ IV.2 Robustness against decoherence
In practical operations, decoherence is inevitable. For the system discussed, we consider two types of noise: single-photon losses and pure dephasing. The master equation of the system can be written as:
@@
where $\rho$ is the density operator of the system, $D[o]\rho=o\rho o^{{\dagger}}-\frac{1}{2}(o^{{\dagger}}o\rho+\rho o^{{\dagger}}o)$. For simplicity, we assume that $\kappa_{k}=\kappa$ $(k=1,2)$ is the single-photon losses rate, and $\gamma_{k}=\gamma$ is the pure dephasing rate. Note that the influence of decoherence in the cavity mode $\mathcal{A}_{0}$ is different from those in the KPOs. The decoherence of the cavity mode $\mathcal{A}_{0}$ can be adiabatically eliminated for large $\Delta$ Chen et al. (2022), so the system is insensitive to the decoherence of $\mathcal{A}_{0}$. In the numerical simulation, we set $\kappa_{0}=\kappa$ and $\gamma_{0}=\gamma$ for convenience. To well understand the effect of decoherence, we project the system onto the eigenstates of $H_{k}^{\mathrm{Kerr}}$. The projection operator of the KPOs is
@@
Figure 7: (a) Fidelities versus the single-photon losses rate $\kappa\in[0,0.1]$ $\mathrm{MHz}$ and the dephasing rate $\gamma=0$. (b) Fidelities versus dephasing $\gamma\in[0,0.1]$ $\mathrm{MHz}$ with single-photon losses rate $\kappa=0$.
Figure 8: Fidelities of the optimized protocol with $N=8$ versus $\kappa=\gamma\in[0,0.1]$ $\mathrm{MHz}$. Other parameters remain unchanged.
@@
When $\gamma_{k}$, $\kappa_{k}$ are much smaller than the energy gap $E_{\mathrm{gap}}$, the dynamics of the system is still well confined to the cat-state subspace Cai et al. (2021); Puri et al. (2019, 2017, 2020). For large $\alpha$, the influence of the single-photon losses in the KPOs is described by the penultimate term in Eq. (50)
@@
$D^{\prime}[a_{k}]\rho\simeq\frac{\alpha^{2}}{\sqrt{1-e^{-4\alpha^{2}}}}D[\sigma^{x}_{k}+ie^{-2\alpha^{2}}\sigma_{k}^{y}]\rho,$ (60)
where $\sigma^{x}_{k}=|C_{+}\rangle_{k}\langle C_{-}|+|C_{-}\rangle_{k}\langle C_{+}|$, $\sigma^{y}_{k}=i(|C_{-}\rangle_{k}\langle C_{+}|+|C_{+}\rangle_{k}\langle C_{-}|)$. Substitute Eq. (60) into Eq. (50), we find that in the computational subspace, single-photon losses mainly lead to a bit-flip error $\sigma^{x}_{k}$, accompanied by an exponentially small phase flip-error $\sigma^{y}_{k}$.
In Fig. 7, we plot the fidelity of the single-photon losses for different values of $N$. The single-photon losses rate is assumed as $\kappa\in[0,0.1]$ $\mathrm{MHZ}$, and the dephasing rate is assumed as $\gamma=0$. As can be seen from Fig. 7, the increase of $N$ does not affect the fidelity too much. The fidelities are all above 0.9 when $\kappa=0.1\mathrm{MHZ}$, which means that the optimized protocol still maintains a high fidelity in the presence of the single-photon losses. We also numerically simulate the dependence of the fidelity on the pure dephasing rate in Fig. 7 and a similar result is obtained.
Since single-photon losses and dephasing may coexist in the system, so we numerically investigate the effect on the fidelity when single-photon losses and dephasing are present together. In Fig. 8, we plot the fidelities of the optimized protocol with $N=8$ versus $\kappa=\gamma\in[0,0.1]$ $\mathrm{MHz}$. As it is shown, when both types of noise are present, the entangled cat states can still maintain high fidelities for small noise rates. However, as $\kappa$ and $\gamma$ gradually increase, the fidelities will be significantly affected Chen et al. (2022).
@@ V CONCLUSION
In conclusion, we have investigated a protocol to use photonic cat-state qubits for preparing high-fidelity entangled cat states. The entangled cat states can be successfully obtained under the action of the cat-code MS gate. We further adopt composite pulses to improve the fidelity of the entangled cat states. The modulated time-independent composite drives can be realized by introducing composite two-photon squeezing drives. Numerical simulation results show that the protocol can provide strong robustness against timing error and detuning error. Furthermore, under the influence of decoherence, the protocol still maintains a high fidelity. That is, stable and high-fidelity entangled cat states are possible with the large $N$. Meanwhile, using composite pulses to enhance the robustness of the protocol is feasible. We hope that the proposal could offer a simple method for preparing stable and high-fidelity entangled cat states in quantum computation.
@@ Acknowledgements.
Y.-H.C. was supported by the National Natural Science Foundation of China under Grant No. 12304390, Fujian 100 Talents Program, and Fujian Minjiang Scholar Program. Y. X. was supported by the National Natural Science Foundation of China under Grant No. 62471143, the Key Program of National Natural Science Foundation of Fujian Province under Grant No. 2024J02008, and the project from Fuzhou University under Grant No. JG2020001-2.
Appendix A Magnus eapansion
[…]
$i\frac{d}{dt}|\psi(t)\rangle=H_{I}(t)|\psi(t)\rangle.$ (61)
The time evolution of the system is determined by the evolution operator $U(t)$
$|\psi(t)\rangle=U(t)|\psi(0)\rangle,$ (62)
where the evolution operator $U(t)$ satisfies the following differential equation
$i\frac{d}{dt}|U(t)\rangle=H_{I}(t)|U(t)\rangle.$ (63)
The Magnus expansion relies on the assumption that there exists an exponential solution, which is given by Chen et al. (2022); Wang et al. (2021)
$U(t)=\exp[-i\sum^{\infty}_{l=1}\Omega_{l}(t)].$ (64)
Under the large detuning condition Wang et al. (2021), only the first two terms of $\sum^{\infty}_{l=1}\Omega_{l}(t)$ are retained. The expressions are as follows
[…]
Using the effective Hamiltonian in Eq. (10) to solve the Eq. (65), we can obtain
$\Omega_{1}(t)$ $=$ $S_{x}(\chi(t)a_{0}^{\dagger}+\chi^{*}(t)a_{0}),$ (69)
$\Omega_{2}(t)$ $=$ $-\beta(t)S^{2}_{x}.$ (71)
Correspondingly, the evolution operator of the system can be derived as
$U(t)=\exp\{-i(\chi(t)a_{0}^{\dagger}+\chi^{*}(t)a_{0})S_{x}+\beta(t)S^{2}_{x}]\},$ (72)
which corresponds to Eq. (11).